# On the Parallel Complexity of Identifying Groups and Quasigroups via Decompositions

### v1

@@ 2.3 Weisfeiler–Leman
$\chi_{k,r+1}(\overline{v})=$ $(\chi_{r}(\overline{v}),\{\!\!\{(\chi_{r}(\overline{v}(v_{1}/x)),\ldots,\chi_{r}(\overline{v}(v_{k}/x))\bigr{|}x\in V(\Gamma)\}\!\!\}),$
@@
$\chi_{k,r+1}(\overline{v})=$ $(\chi_{r}(\overline{v}),\{(\chi_{r}(\overline{v}(v_{1}/x)),\ldots,\chi_{r}(\overline{v}(v_{k}/x))\bigr{|}x\in V(\Gamma)\}).$
@@ 2.4 Pebbling Games
    Spoiler places $p_{i}$ on some vertex $g\in G$. Then $p_{i}^{\prime}$ is placed on $f(g)$.
Suppose that $(g_{1},\ldots,g_{\ell})\mapsto(h_{1},\ldots,h_{\ell})$ have been pebbled. In Version I, Duplicator wins at the given round if this map satisfies the initial coloring condition of WL Version I: (a) for all $i,j,m\in[\ell]$, $g_{i}g_{j}=g_{m}\iff h_{i}h_{j}=h_{m}$, and (b) for all $i,j\in[\ell]$, $g_{i}=g_{j}\iff h_{i}=h_{j}$. In Version II, Duplicator wins at the given round if the map $(g_{1},\ldots,g_{\ell})\mapsto(h_{1},\ldots,h_{\ell})$ extends to an isomorphism of the generated subgroups $\langle g_{1},\ldots,g_{\ell}\rangle$ and $\langle h_{1},\ldots,h_{\ell}\rangle.$ Brachter & Schweitzer established that for $J\in\{I,II\}$, $(k,r)$-WL Version J is equivalent to version J of the $(k+1)$-pebble, $r$-round pebble game [BS20].
@@
There exist analogous pebble games for count-free WL Versions I-II. The count-free $(k+1)$-pebble game consists of two players: Spoiler and Duplicator, as well as $(k+1)$ pebble pairs $(p,p^{\prime})$. In both versions, Spoiler wishes to show that the two groups $G$ and $H$ are not isomorphic. Duplicator wishes to show that the two groups are isomorphic. Each round of the game proceeds as follows.
@@
Spoiler wins, by definition, at round 0 if $G$ and $H$ do not have the same number of elements. Suppose that $|G|=|H|$, and let $\bar{g}\in G^{k},\bar{h}\in H^{k}$. We have that $\bar{g},\bar{h}$ are not distinguished by the first $r$ rounds of the count-free $k$-WL version $J$ if and only if Duplicator wins the first $r$ rounds of the count-free Version $J$ $(k+1)$-pebble game [IL90, CFI92, GL23b].
@@ 2.5 Weisfeiler–Leman as a Parallel Algorithm
  • WL Version I: Let $(g_{1},\ldots,g_{k})$ and $(h_{1},\ldots,h_{k})$ be two $k$-tuples of group elements. We may test in $\textsf{AC}^{0}$ whether (a) for all $i,j,m\in[k]$, $g_{i}g_{j}=g_{m}\iff h_{i}h_{j}=h_{m}$, and (b) $g_{i}=g_{j}\iff h_{i}=h_{j}$. So we may decide if two $k$-tuples receive the same initial color in $\textsf{AC}^{0}$. Comparing the (multi)set of colors at the end of each iteration (including after the initial coloring), as well as the refinement steps, proceed identically as in [GV06]. Thus, for fixed $k$, each iteration of the counting $k$-WL Version I can be implemented using a DLOGTIME-uniform $\textsf{TC}^{0}$ circuit, and each iteration of the count-free $k$-WL Version I can be implemented using a DLOGTIME-uniform $\textsf{AC}^{0}$ circuit.
@@ 3.1 Additional Preliminaries for Direct Products
Let $G$ be a group, and let $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ be a fully refined direct product decomposition of $G$. For any $i\in[k]$, we have that $S_{i}$ appears in any fully refined direct product decomposition of $G$ if and only if for all distinct $j\neq i$, $|S_{i}/[S_{i},S_{i}]|$ and $|Z(G_{j})|$ are coprime. In particular, $\mathcal{S}$ is unique if and only if for all distinct $i,j\in[k]$, $|S_{i}/[S_{i},S_{i}]|$ and $|Z(G_{j})|$ are coprime.
@@
Conversely, suppose that there is a $j\neq i$ such that $|S_{i}/[S_{i},S_{i}]|$ and $|Z(S_{j})|$ are not coprime. We will construct a fully refined direct product decomposition of $G$ in which $S_{i}$ does not appear. Let $\beta_{ij}:S_{i}\to Z(S_{j})$ be any nontrivial homomorphism. Set $\beta_{iv}=1$ for every $v\neq j$, $\beta_{i}=\prod_{v=1}^{k}\beta_{iv}$, $\beta_{u}=1$ for every $u\neq i$, and $\beta=\prod_{u=1}^{k}\beta_{u}$. Then $\beta$ is a homomorphism $G\to Z(G)$. Let $\alpha:G\to Z(G)$ be given by $\alpha(g)=g\beta(g)$. Since $\beta(G)\subseteq Z(G)$, $\alpha$ is a central homomorphism. We will show that $\alpha$ is an automorphism. It suffices to show that $\alpha$ is injective. Let $g=(g_{1},\dots,g_{k})\in G=\prod_{u=1}^{k}S_{u}$. Since $\beta_{ij}$ is the only nontrivial component of $\beta$, we have $\alpha(g_{1},\dots,g_{k})=(g_{1},\dots,g_{j-1},g_{j}^{\prime},g_{j+1},\dots,g_{k})$, where $g_{j}^{\prime}=g_{j}\beta_{ij}(g_{i})$. Suppose that $h=(h_{1},\dots,h_{k})\in G$ and $\alpha(g)=\alpha(h)$. Then $g_{v}=h_{v}$ for every $v\neq j$, and $g_{j}\beta_{ij}(g_{i})=h_{j}\beta_{ij}(h_{i})$. As $g_{i}=h_{i}$, we conclude that $g_{j}=h_{j}$, too, and $g=h$.
@@ 3.2 Weisfeiler–Leman and Direct Products
Let $G\in\mathcal{C}$ (see Definition 1.1) be a group of order $n$. Our first step will be to show that if $H\not\in\mathcal{C}$, then count-free WL will distinguish $H$ from in $O(\log\log n)$ rounds. We will accomplish this using the Rank Lemma due to Grochow and Levet [GL23b, Lemma 4.3]. We will first introduce a notion of weight.
Definition 3.7 (cf. [GL23a, Definition 4.9]).
@@
Definition 3.7 generalizes [GL23a, Definition 4.9], where Grochow and Levet considered $G$ to have no Abelian normal subgroups and $H$ to be the socle of $G$, denoted $\text{Soc}(G)$. In this setting, the $S_{i}$ in the unique refined direct product decomposition of $\text{Soc}(G)$ were the non-Abelian simple direct factors of $\text{Soc}(G)$.
As in [GL23a], our notion of weight is a special case of rank, as introduced in [GL23b].
Definition 3.9 ([GL23b, Definition 4.1]).
@@
Lemma 3.10 ([GL23b, Rank Lemma 4.3]).
@@
Let $G$ be a group of order $n$, and let $S\trianglelefteq G$ be $O(1)$-generated. Let $H$ be an arbitrary group. Let $g\in S$, $h\in H$. Suppose that there does not exist $T\trianglelefteq H$ such that $h\in T$ and $T\cong S$. Then Spoiler can win with $O(1)$ pebbles and $O(1)$ rounds, from the initial configuration $g\mapsto h$.
@@
  • Case 1: Suppose first that there does not exist some $T\leq H$ such that $h\in T$ and $T\cong S$. Spoiler uses $k$ pebbles and $k$ rounds to pebble generators $(g_{1},\ldots,g_{k})$ for $S$. Let $(h_{1},\ldots,h_{k})$ be Duplicator’s response. As $H$ does not contain a copy of $S$, it follows that $(g,g_{1},\ldots,g_{k})\mapsto(h,h_{1},\ldots,h_{k})$ does not extend to an isomorphism of $S=\langle g,g_{1},\ldots,g_{k}\rangle$ and $\langle h,h_{1},\ldots,h_{k}\rangle$. So Spoiler wins with $k$ pebbles and $k$ rounds.
  • Case 2: Now suppose instead that there exists some $T\leq H$ such that $h\in T$ and $T\cong S$. Suppose however, that for all such $T$, we have that $T\not\trianglelefteq H$. Spoiler begins by pebbling generators $(g_{1},\ldots,g_{k})$ for $S$. Let $(h_{1},\ldots,h_{k})$ be Duplicator’s response. We may assume that the map $(g,g_{1},\ldots,g_{k})\mapsto(h,h_{1},\ldots,h_{k})$ extends to an isomorphism of $S=\langle g_{1},\ldots,g_{k}\rangle$ and $T:=\langle h_{1},\ldots,h_{k}\rangle$. Otherwise, Spoiler wins. As $T\not\trianglelefteq H$, there exists some $x\in H$ such that $xTx^{-1}\neq T$. Spoiler pebbles such an $x\in H$, and Duplicator responds by pebbling some $y\in G$. As $S$ is normal in $G$, it follows that $ySy^{-1}=S$. Thus, the map $(g,g_{1},\ldots,g_{k},y)\mapsto(h,h_{1},\ldots,h_{k},x)$ does not extend to an isomorphism of $\langle g,g_{1},\ldots,g_{k},y\rangle$ and $\langle h,h_{1},\ldots,h_{k},x\rangle$. So Spoiler wins with $k+1$ pebbles and in $k+1$ rounds. ∎
@@
Thus, Spoiler now reuses the pebbles on $g,g^{\prime\prime}$, and we iterate on this strategy starting from $g^{\prime}\mapsto h^{\prime}$. Thus, after $O(\log k)\leq O(\log\log n)$ rounds, we have reached the case where $\text{wt}(g^{\prime})=\text{wt}(h^{\prime})=1$. Both cases are handled by Lemma 3.11; thus, Spoiler now wins with $O(1)$ additional pebbles and $O(1)$ additional rounds. In total, Spoiler used $O(1)$ additional pebbles and $O(\log\log n)$ additional rounds. ∎
@@
Let $H$ be a group of order $n:=|G|$ such that $H\not\cong G$. Duplicator chooses a bijection $f:G\to H$. As $G\not\cong H$, there exists some $g\in G$ such that $\text{wt}(g)=1$, and either: (i) $\text{wt}(f(g))>1$, or (ii) $\text{wt}(g)=1$ and the (unique) direct factor $S\trianglelefteq G$ containing $g$ is not isomorphic to the (unique) direct factor $T\trianglelefteq H$ containing $f(g)$. By Lemma 3.11, Spoiler wins with $O(1)$ pebbles and $O(1)$ rounds. The result now follows. ∎
@@
The Classification of Finite Simple Groups provides that every finite simple group is generated by at most two elements. However, there are no bounds in general of the length of the words for such generating sets.
@@ 4 Direct Product Decompositions in Parallel
In this section, we will prove Theorem 1.7.
@@ 4.1 Group Division in Parallel
Let $G$ be a group, and let $x\in G$. We can compute $|x|$ in $\textsf{L}\cap\textsf{FOLL}$.
@@
The complexity of order finding was recently improved to $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\textsf{L}\cap\mathsf{FOLL}$ [CGLW24].
@@
We first decompose $G$ into a direct product of its Sylow subgroups. As $G$ is Abelian, we have that for each prime $p\mid|G|$, there exists exactly one Sylow $p$-subgroup $A_{p}$ of $G$. We may, in $\textsf{L}\cap\textsf{FOLL}$, compute the Sylow subgroups of $G$ [BKLM01, Lemma 3.1]. It remains to directly decompose each Sylow subgroup of $G$, which we will do in parallel.
Fix a prime $p$, and consider the Sylow $p$-subgroup $A_{p}$ of $G$. We proceed as follows. First, for each $g\in A_{p}$, we test whether $g$ splits from $A_{p}$. For some fixed $g$, we may use Lemma 4.3 to decide in $\textsf{L}\cap\textsf{FOLL}$ whether there exists some $h\in A_{p}$ such that $|gh^{p}|<|g|$. By Lemma 4.2, we have that $g$ splits from $A_{p}$ precisely if no such $h$ exists. Thus, in $\textsf{L}\cap\textsf{FOLL}$, we can identify the elements of $A_{p}$ that split from $A_{p}$. Next, using Lemma 4.3, we may for each $g\in A_{p}$, write down the elements of $\langle g\rangle$.
@@
Note that only $\log|A_{p}|\leq\log|G|$ many iterations are required to compute a fully-refined direct product decomposition of $A_{p}$. Our pre-processing steps in the first two paragraphs are L-computable, and the remaining work is $\textsf{AC}^{1}$-computable. Thus, we may compute a fully-refined direct product decomposition of $A_{p}$ in $\textsf{AC}^{1}$. The result now follows. ∎
@@
  • Case 1: Suppose that $G/A$ is Abelian. Given $g,h\in G$, we can decide in $\textsf{AC}^{0}$ whether $g$ and $h$ belong to the same coset of $G/A$, by checking whether $gh^{-1}\in A$. Thus, we may write down the multiplication table for $G/A$ in $\textsf{AC}^{0}$. As $G/A$ is Abelian, we have by Proposition 4.5 that we have that we can compute a fully-refined direct product decomposition for $G/A=\langle g_{1}A\rangle\times...\times\langle g_{k}A\rangle$ in $\textsf{AC}^{1}$. Now, for each coset $g_{i}A$ ($i\in[k]$), it remains to select a representative $b_{i}\in g_{i}A$ such that $b_{i}\in Z(G)$ and $|b_{i}|=|g_{i}A|$. By Lemma 4.3, this step is computable in $\textsf{L}\cap\mathsf{FOLL}$. The total work in this case is $\textsf{AC}^{1}$, as desired.
  • Case 2: Suppose instead that $G/A$ is not Abelian. In $\mathsf{AC^{0}}$, we may decide whether two elements of $G$ commute. Thus, in $\mathsf{AC^{0}}$, we may write down the elements of $C_{G}(A)$. We now turn to computing $T=[C_{G}(A),G]$. In $\mathsf{AC^{0}}$, we may write down $S:=\{[a,g]:a\in C_{G}(A),g\in G\}$. Now using a membership test, we may in L, write down $T=\langle S\rangle$ [BM91, Rei08].
@@
    For each $i\in[k]$, we select a coset representative $c_{i}\in b_{i}T$ (from Kayal and Nezhmetdinov [KN09], we have that any such coset representative suffices). Now in L, we may write down $C:=\langle T,c_{1},\ldots,c_{k}\rangle$ using a membership test [BM91, Rei08]. We may now check in $\mathsf{AC^{0}}$ whether $G=A\times C$. If $G=A\times C$, then we return $C$. Otherwise, we return that no such decomposition exists.

### v2

@@ 2.3 Weisfeiler–Leman
$\chi_{k,r+1}(\overline{v})=$ $(\chi_{r}(\overline{v}),\{\!\!\{(\chi_{r}(\overline{v}(v_{1}/x)),\ldots,\chi_{r}(\overline{v}(v_{k}/x)))\bigr|x\in V(\Gamma)\}\!\!\}),$
@@
$\chi_{k,r+1}(\overline{v})=$ $(\chi_{r}(\overline{v}),\{(\chi_{r}(\overline{v}(v_{1}/x)),\ldots,\chi_{r}(\overline{v}(v_{k}/x)))\bigr|x\in V(\Gamma)\}).$
@@
The terms distinguish and identify also extend in the natural way when controlling only for the dimension $k$. The (count-free) Weisfeiler–Leman dimension (WL-dimension) of a graph $G$ is the minimum $k$, such that $k$-WL identifies $G$.
@@ 2.4 Pebbling Games
    Spoiler places $p_{i}$ on some vertex $g\in G$. Then $p_{i}^{\prime}$ is placed on $f_{r}(g)$.
Suppose that $(g_{1},\ldots,g_{\ell})\mapsto(h_{1},\ldots,h_{\ell})$ have been pebbled. In Version I, Duplicator wins at the given round if this map satisfies the initial coloring condition of WL Version I: (a) for all $i,j,m\in[\ell]$, $g_{i}g_{j}=g_{m}\iff h_{i}h_{j}=h_{m}$, and (b) for all $i,j\in[\ell]$, $g_{i}=g_{j}\iff h_{i}=h_{j}$. In Version II, Duplicator wins at the given round if the map $(g_{1},\ldots,g_{\ell})\mapsto(h_{1},\ldots,h_{\ell})$ extends to an isomorphism of the generated subgroups $\langle g_{1},\ldots,g_{\ell}\rangle$ and $\langle h_{1},\ldots,h_{\ell}\rangle.$ Brachter & Schweitzer established that for $J\in\{I,II\}$, two $k$-tuples $\bar{g}\in G^{k},\bar{h}\in H^{k}$ receive different colors under $(k,r)$-WL Version $J$ if and only if Spoiler has a winning strategy in the $(k+1)$-pebble, $r$-round pebble game starting from the initial configuration $\bar{g}\mapsto\bar{h}$ [BS20].
@@
There exist analogous pebble games for count-free WL Versions I-II. The count-free $(k+1)$-pebble game consists of two players: Spoiler and Duplicator, as well as $(k+1)$ pebble pairs $(p,p^{\prime})$. In both versions, Spoiler wishes to show that the two groups $G$ and $H$ are not isomorphic. Duplicator wishes to show that the two groups are isomorphic. Prior to the start of the game, each pebble pair $(p_{i},p_{i}^{\prime})$ is initially placed either beside the groups or on a given pair of group elements $v_{i}\mapsto v_{i}^{\prime}$ (where $v_{i}\in G,v_{i}^{\prime}\in H$). Each round of the game proceeds as follows.
@@
Spoiler wins, by definition, at round 0 if $G$ and $H$ do not have the same number of elements. Suppose that $|G|=|H|$, and let $\bar{g}\in G^{k},\bar{h}\in H^{k}$. We have that for $J\in\{I,II\}$, $\bar{g}$ and $\bar{h}$ receive different colors under the count-free $(k,r)$-WL Version $J$ if and only if Spoiler has a winning strategy in the $(k+1)$-pebble, $r$-round count-free pebble game starting from the initial configuration $\bar{g}\mapsto\bar{h}$ [IL90, CFI92, GL26].
@@ 2.5 Weisfeiler–Leman as a Parallel Algorithm
  • WL Version I: Let $(g_{1},\ldots,g_{k})$ and $(h_{1},\ldots,h_{k})$ be two $k$-tuples of group elements. We may test in $\textsf{AC}^{0}$ whether (a) for all $i,j,m\in[k]$, $g_{i}g_{j}=g_{m}\iff h_{i}h_{j}=h_{m}$, and (b) $g_{i}=g_{j}\iff h_{i}=h_{j}$. So we may decide if two $k$-tuples receive the same initial color in $\textsf{AC}^{0}$. Comparing the (multi)set of colors at the end of each iteration (including after the initial coloring), as well as the refinement steps, proceeds identically as in [GV06]. Thus, for fixed $k$, each iteration of the counting $k$-WL Version I can be implemented using a DLOGTIME-uniform $\textsf{TC}^{0}$ circuit, and each iteration of the count-free $k$-WL Version I can be implemented using a DLOGTIME-uniform $\textsf{AC}^{0}$ circuit.
@@ 3.1 Additional Preliminaries for Direct Products
Let $G$ be a group, and let $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ be a fully refined direct product decomposition of $G$. For any $i\in[k]$, we have that $S_{i}$ appears in any fully refined direct product decomposition of $G$ if and only if for all distinct $j\neq i$, $|S_{i}/[S_{i},S_{i}]|$ and $|Z(S_{j})|$ are coprime. In particular, $\mathcal{S}$ is unique if and only if for all distinct $i,j\in[k]$, $|S_{i}/[S_{i},S_{i}]|$ and $|Z(S_{j})|$ are coprime.
@@
Conversely, suppose that there is a $j\neq i$ such that $|S_{i}/[S_{i},S_{i}]|$ and $|Z(S_{j})|$ are not coprime. We will construct a fully refined direct product decomposition of $G$ in which $S_{i}$ does not appear. Let $\beta_{ij}:S_{i}\to Z(S_{j})$ be any nontrivial homomorphism. Set $\beta_{iv}=1$ for every $v\neq j$, $\beta_{i}=\prod_{v=1}^{k}\beta_{iv}$, $\beta_{u}=1$ for every $u\neq i$, and $\beta=\prod_{u=1}^{k}\beta_{u}$. Then $\beta$ is a homomorphism $G\to Z(G)$. Let $\alpha:G\to G$ be given by $\alpha(g)=g\beta(g)$. Since $\beta(G)\subseteq Z(G)$, $\alpha$ is a central homomorphism. We will show that $\alpha$ is an automorphism. It suffices to show that $\alpha$ is injective. Let $g=(g_{1},\dots,g_{k})\in G=\prod_{u=1}^{k}S_{u}$. Since $\beta_{ij}$ is the only nontrivial component of $\beta$, we have $\alpha(g_{1},\dots,g_{k})=(g_{1},\dots,g_{j-1},g_{j}^{\prime},g_{j+1},\dots,g_{k})$, where $g_{j}^{\prime}=g_{j}\beta_{ij}(g_{i})$. Suppose that $h=(h_{1},\dots,h_{k})\in G$ and $\alpha(g)=\alpha(h)$. Then $g_{v}=h_{v}$ for every $v\neq j$, and $g_{j}\beta_{ij}(g_{i})=h_{j}\beta_{ij}(h_{i})$. As $g_{i}=h_{i}$, we conclude that $g_{j}=h_{j}$, too, and $g=h$.
@@ 3.2 Weisfeiler–Leman and Direct Products
Let $G\in\mathcal{C}$ (see Definition 1.1) be a group of order $n$. Our first step will be to show that if $H\not\in\mathcal{C}$, then count-free WL Version II will distinguish $H$ from in $O(\log\log n)$ rounds. We will accomplish this using the Rank Lemma due to Grochow and Levet [GL26, Lemma 4.3]. We will first introduce a notion of weight.
Definition 3.7 (cf. [GL25, Definition 4.9]).
@@
Definition 3.7 generalizes [GL25, Definition 4.9], where Grochow and Levet considered $G$ to have no Abelian normal subgroups and $H$ to be the socle of $G$, denoted $\text{Soc}(G)$. (Note that $\text{Soc}(G)$ is the direct product of the minimal normal subgroups of $G$.) In this setting, the $S_{i}$ in the unique refined direct product decomposition of $\text{Soc}(G)$ were the non-Abelian simple direct factors of $\text{Soc}(G)$.
As in [GL25], our notion of weight is a special case of rank, as introduced in [GL26].
Definition 3.9 ([GL26, Definition 4.1]).
@@
Lemma 3.10 ([GL26, Rank Lemma 4.3]).
@@
Let $G$ be a group of order $n$, and let $S\trianglelefteq G$ be $O(1)$-generated. Let $H$ be an arbitrary group. Let $g\in S$, $h\in H$. Suppose that there does not exist $T\trianglelefteq H$ such that $h\in T$ and $T\cong S$. Then in the Version II pebble game, Spoiler can win with $O(1)$ pebbles and $O(1)$ rounds, from the initial configuration $g\mapsto h$.
@@
  • Case 1: Suppose first that there does not exist some $T\leq H$ such that $h\in T$ and $T\cong S$. Spoiler uses $k$ pebbles and $k$ rounds to pebble generators $(g_{1},\ldots,g_{k})$ for $S$. Let $(h_{1},\ldots,h_{k})$ be Duplicator’s response. As there does not exist a copy of $S$ in $H$ that also contains $h$, it follows that $(g,g_{1},\ldots,g_{k})\mapsto(h,h_{1},\ldots,h_{k})$ does not extend to an isomorphism of $S=\langle g,g_{1},\ldots,g_{k}\rangle$ and $\langle h,h_{1},\ldots,h_{k}\rangle$. So Spoiler wins with $k$ pebbles and $k$ rounds.
  • Case 2: Now suppose instead that there exists some $T\leq H$ such that $h\in T$ and $T\cong S$. Suppose however, that for all such $T$, we have that $T\not\trianglelefteq H$. Spoiler begins by pebbling generators $(g_{1},\ldots,g_{k})$ for $S$. Let $(h_{1},\ldots,h_{k})$ be Duplicator’s response. We may assume that the map $(g,g_{1},\ldots,g_{k})\mapsto(h,h_{1},\ldots,h_{k})$ extends to an isomorphism of $S=\langle g_{1},\ldots,g_{k}\rangle$ and $T:=\langle h_{1},\ldots,h_{k}\rangle$ (and in particular, that $h\in T$). Otherwise, Spoiler wins. As $T\not\trianglelefteq H$, there exists some $x\in H$ such that $xTx^{-1}\neq T$. Spoiler pebbles such an $x\in H$, and Duplicator responds by pebbling some $y\in G$. As $S$ is normal in $G$, it follows that $ySy^{-1}=S$. Thus, the map $(g,g_{1},\ldots,g_{k},y)\mapsto(h,h_{1},\ldots,h_{k},x)$ does not extend to an isomorphism of $\langle g,g_{1},\ldots,g_{k},y\rangle$ and $\langle h,h_{1},\ldots,h_{k},x\rangle$. So Spoiler wins with $k+1$ pebbles and in $k+1$ rounds. ∎
@@
Thus, Spoiler now reuses the pebbles on $g,g^{\prime\prime}$, and we iterate on this strategy starting from $g^{\prime}\mapsto h^{\prime}$. Thus, after $O(\log k)\leq O(\log\log n)$ rounds, we have reached the case where $\text{wt}(g^{\prime})=\text{wt}(h^{\prime})=1$. By Lemma 3.11, Spoiler now wins with $O(1)$ additional pebbles and $O(1)$ additional rounds. In total, Spoiler used $O(1)$ pebbles and $O(\log\log n)$ rounds. ∎
@@
Let $H$ be a group of order $n:=|G|$ such that $H\not\cong G$. Duplicator chooses a bijection $f:G\to H$. As $G\not\cong H$, there exists some $g\in G$ such that $\text{wt}(g)=1$, and either: (i) $\text{wt}(f(g))>1$, or (ii) $\text{wt}(g)=1$ and the (unique) direct factor $S\trianglelefteq G$ containing $g$ is not isomorphic to the (unique) direct factor $T\trianglelefteq H$ containing $f(g)$. Both of these cases are handled by Lemma 3.11, which provides that Spoiler wins with $O(1)$ pebbles and $O(1)$ rounds. The result now follows. ∎
@@
The Classification of Finite Simple Groups provides that every finite simple group is generated by at most two elements $\{a,b\}$. As the diameter of the Cayley graph $\text{Cay}(S,\{a,b\})$ is at most $|S|$, it follows that every element in $S$ can be realized as a word of length at most $|S|$ over $\{a,b\}$. We are not aware of better bounds for such $2$-element generating sets.
@@ 4 Direct Product Decompositions in Parallel
In this section, we will prove Theorem 1.7. We accomplish this by parallelizing the work of Kayal and Nezhmetdinov [KN09].
@@ 4.1 Group Division in Parallel
Let $G$ be a group, and let $x\in G$. We can compute $|x|$, as well as $\langle x\rangle$, in $\textsf{FL}\cap\textsf{FOLL}$.
@@
The complexity of order finding was recently improved to $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\textsf{FL}\cap\mathsf{FOLL}$ [CGLW25].
@@
It is possible to compute a basis for an Abelian group in $\textsf{NC}^{2}$ using Smith normal form [Vil94]. Proposition 4.5 improves this bound to $\textsf{AC}^{1}$, when the underlying group is given by its multiplication table.
@@
We first decompose $G$ into a direct product of its Sylow subgroups. As $G$ is Abelian, we have that for each prime $p\mid|G|$, there exists exactly one Sylow $p$-subgroup $A_{p}$ of $G$. We may, in $\textsf{FL}\cap\textsf{FOLL}$, compute the Sylow subgroups of $G$ [BKLM01, Lemma 3.1]. It remains to directly decompose each Sylow subgroup of $G$, which we will do in parallel.
Fix a prime $p$, and consider the Sylow $p$-subgroup $A_{p}$ of $G$. We proceed as follows. First, for each $g\in A_{p}$, we test whether $g$ splits from $A_{p}$. For some fixed $g$, we may use Lemma 4.3 to decide in $\textsf{FL}\cap\textsf{FOLL}$ whether there exists some $h\in A_{p}$ such that $|gh^{p}|<|g|$. By Lemma 4.2, we have that $g$ splits from $A_{p}$ precisely if no such $h$ exists. Thus, in $\textsf{FL}\cap\textsf{FOLL}$, we can identify the elements of $A_{p}$ that split from $A_{p}$. Next, using Lemma 4.3, we may for each $g\in A_{p}$, write down the elements of $\langle g\rangle$ in $\textsf{FL}\cap\textsf{FOLL}$.
@@
Note that only $\log|A_{p}|\leq\log|G|$ many iterations are required to compute a fully-refined direct product decomposition of $A_{p}$. Our pre-processing steps in the first two paragraphs are FL-computable, and the remaining work is $\textsf{AC}^{1}$-computable. Thus, we may compute a fully-refined direct product decomposition of $A_{p}$ in $\textsf{AC}^{1}$. The result now follows. ∎
@@
  • Case 1: Suppose that $G/A$ is Abelian. Given $g,h\in G$, we can decide in $\textsf{AC}^{0}$ whether $g$ and $h$ belong to the same coset of $G/A$, by checking whether $gh^{-1}\in A$. Thus, we may write down the multiplication table for $G/A$ in $\textsf{AC}^{0}$. As $G/A$ is Abelian, we have by Proposition 4.5 that we have that we can compute a fully-refined direct product decomposition for $G/A=\langle g_{1}A\rangle\times...\times\langle g_{k}A\rangle$ in $\textsf{AC}^{1}$. Now, for each coset $g_{i}A$ ($i\in[k]$), it remains to select a representative $b_{i}\in g_{i}A$ such that $b_{i}\in Z(G)$ and $|b_{i}|=|g_{i}A|$. By Lemma 4.3, this step is computable in $\textsf{FL}\cap\mathsf{FOLL}$. The total work in this case is $\textsf{AC}^{1}$-computable, as desired.
  • Case 2: Suppose instead that $G/A$ is not Abelian. In $\mathsf{AC^{0}}$, we may decide whether two elements of $G$ commute. Thus, in $\mathsf{AC^{0}}$, we may write down the elements of $C_{G}(A)$. We now turn to computing $T=[C_{G}(A),G]$. In $\mathsf{AC^{0}}$, we may write down $S:=\{[a,g]:a\in C_{G}(A),g\in G\}$. Now using a membership test, we may in FL, write down $T=\langle S\rangle$ [BM91, Rei08].
@@
    For each $i\in[k]$, we select a coset representative $c_{i}\in b_{i}T$; from Kayal and Nezhmetdinov [KN09], we have that any such coset representative suffices. Now in FL, we may write down $C:=\langle T,c_{1},\ldots,c_{k}\rangle$ using a membership test [BM91, Rei08]. We may now check in $\mathsf{AC^{0}}$ whether $G=A\times C$. If $G=A\times C$, then we return $C$. Otherwise, we return that no such decomposition exists. The total work in this case is $\textsf{AC}^{1}$-computable, as desired.